OpenStudy (anonymous):

Hello. How can I solve cos(γ+135°)+cos(γ-135°)?

myininaya (myininaya):

try using the sum and difference identities for cosine also i don't think the instructions say solve

myininaya (myininaya):

\[\cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b) \\ \cos(a-b)=\cos(a)\cos(b)+\sin(a)\sin(b)\]

OpenStudy (anonymous):

Everything I was able to do with it was this: cos(γ+135°)+cos(γ-135°) = cosγ*cos135°-sinγ*sin135°+cosγ*cos135°+sinγ*sin135° =cosγ*cos135°*sin²135° and the book says I'm supposed to get this: -√2cosγ

OpenStudy (anonymous):

cos(a+b) = cos(a)cos(b) - sin(a)sin(b) cos(a-b) = cos(a)cos(b) + sin(a)sin(b) cos(y+135)+cos(y-135) = (cosy*cos135- siny*sin135)+(cosy*cos135+siny*sin135) Put value of cos135 and sin135 or you can further expand it by making 135 = 90+45

OpenStudy (anonymous):

I have given the solution, do check it

OpenStudy (shinalcantara):

\[\cos( \gamma + 135^o) + \cos( \gamma - 135^o)\] \[(\cos \gamma \cos 135^o - \sin \gamma \sin 135^o) + (\cos \gamma \cos 135^o +\sin \gamma \sin135^o)\] removing the parenthesis: \[\cos \gamma \cos 135^o - \sin \gamma \sin 135^o + \cos \gamma \cos 135^o+\sin \gamma \sin135^o\] \[\sin \gamma \sin 135^o - \sin \gamma \sin 135^o \rightarrow 0\] \[\cos \gamma \cos 135^o + \cos \gamma \cos 135^o\] \[2\cos \gamma \cos 135^o\] \[\cos 135^o \rightarrow \frac{ -\sqrt{2} }{ 2 }\] \[2 \frac{ -\sqrt{2} }{ 2 } \cos \gamma \rightarrow -\sqrt{2}\cos \gamma\]

OpenStudy (anonymous):

thank you (: